# Token-Efficient Long Video Understanding for Multimodal LLMs

### v1

@@ Abstract
Recent advances in video-based multimodal large language models (Video-LLMs) have significantly improved video understanding by processing videos as sequences of image frames. However, many existing methods treat frames independently in the vision backbone, lacking explicit temporal modeling, which limits their ability to capture dynamic patterns and efficiently handle long videos. To address these limitations, we introduce STORM (Spatiotemporal TOken Reduction for Multimodal LLMs), a novel architecture incorporating a dedicated temporal encoder between the image encoder and the LLM. Our temporal encoder leverages the Mamba State Space Model to integrate temporal information into image tokens, generating enriched representations that preserve inter-frame dynamics across the entire video sequence. This enriched encoding not only enhances video reasoning capabilities but also enables effective token reduction strategies, including test-time sampling and training-based temporal and spatial pooling, substantially reducing computational demands on the LLM without sacrificing key temporal information. By integrating these techniques, our approach simultaneously reduces training and inference latency while improving performance, enabling efficient and robust video understanding over extended temporal contexts. Extensive evaluations show that STORM achieves state-of-the-art results across various long video understanding benchmarks (more than 5% improvement on MLVU and LongVideoBench) while reducing the computation costs by up to and the decoding latency by 2.4-2.9 for the fixed numbers of input frames.
@@ 1 Introduction
Recent advancements in video-based multimodal Large Language Models (Video-LLMs) have significantly improved the ability of artificial intelligence systems to understand and generate descriptions of video content [1, 2, 3, 4, 5, 6]. A common strategy in these models involves treating a video as a sequence of individual image frames, processing each frame independently using an image encoder and a visual-language projector [3, 4, 2]. The resulting frame-level representations are then fed into a large language model (LLM), which performs temporal reasoning to comprehend the sequence of events depicted in the video. Despite leveraging powerful image encoders and language understanding capabilities of LLMs, this approach exhibits fundamental limitations in video processing, particularly for long videos. The absence of explicit temporal encoding fails to capture crucial temporal dynamics between frames, forcing the LLM to infer temporal relationships solely from sequences of static images. This sequential processing imposes a substantial computational burden on the language model, significantly impairing its ability to handle extended video sequences and the generalization ability to longer contexts during inference. To manage computational overhead, existing methods often rely on naive frame subsampling to reduce the number of tokens for LLM processing. However, this approach results in significant information loss from discarded frames, potentially eliminating critical details necessary for comprehensive video understanding [7, 8]. Additionally, these methods fail to effectively compress consecutive frames that typically contain substantial overlapping information. 000Website: https://research.nvidia.com/labs/lpr/storm
In this work, we propose STORM (Spatiotemporal TOken Reduction for Multimodal LLMs), which introduces a novel temporal encoder between the image encoder and the LLM that bridges the gap between visual and language representations. This architecture integrates temporal dynamics earlier in the pipeline, significantly enhancing the temporal reasoning capabilities of Video-LLMs while enabling substantial downstream computational efficiency. By injecting temporal information directly into visual tokens, we reduce the LLM’s temporal reasoning burden, allowing it to focus on higher-level language tasks. We employ the Mamba State Space Model [9] as the core of our temporal layer, enabling efficient processing of long videos while enhancing generalization to extended temporal contexts. The temporal layer processes image and video inputs differently — for images, it functions as a spatial scanner that enhances tokens by incorporating global spatial context; for videos, it performs simultaneous spatial and temporal scanning that captures comprehensive spatiotemporal information.
@@
The key advantage of the Mamba layer is its ability to compress historical information into state representations. As consecutive frames in the video input often contain redundant information, our temporal encoder efficiently processes and propagates temporal information across the entire video sequence. The resulting visual tokens inherently encapsulate temporal history, effectively summarizing the temporal dynamics of the video. This characteristic allows us to extract fewer visual tokens for LLM processing while preserving key information. In our experiments, we explore a straightforward yet effective token reduction approach by applying training-free token subsampling at test time. This method not only reduces computational costs but also enhances performance across various scenarios. Additionally, we also explore training-based compression method including temporal and spatial token compression. Temporal pooling reduces the number of tokens along the temporal dimension, while spatial average pooling decreases tokens per frame. These compression strategies are optimized during training to preserve essential information while minimizing redundancy. Unlike previous methods that directly subsample raw video frames, which can lead to loss of critical information, our approaches preserve the essential temporal information in a compressed format. Our method not only reduces the computational load on the LLM but also improves model performance through more comprehensive video representation in a compact token space. Our empirical evaluation shows that the proposed approach successfully extends to the long-context video understanding model without compromising training efficiency.
@@ 3 Method
This section presents the Mamba-based temporal projector architecture and introduces several token compression techniques for efficient long video processing. We propose two token compression approaches: temporal and spatial compression. We first detail our temporal and spatial pooling strategies that effectively reduce the number of tokens during training. Additionally, we propose a test-time temporal token sampling method that maintains model performance without requiring additional training steps. The overview of our method is shown in Figure 2.
@@ 4 Experiments
4.1 Training Details
We utilize the pre-trained SigLIP [12] from PaliGemma [27] and Qwen2-VL [28], respectively, and fine-tune them to adapt to our video datasets. The temporal projector is initialized with random weights. Each image is always resized to a resolution.
@@ 4.1 Training Details
VILA-Based Models Size Comp. Latency #Frames # Frames MVBench MLVU LongVideoBench VideoMME Ratio (%) (s) (train) (test) test dev val (w/o sub.) Duration 16 sec 3120 min 8 sec60 min 160 min Token Budget: 8K VILA Baseline 7B 100 4.38 32 256 69.5 70.2 55.9 60.1 STORM 7B 100 4.26 32 256 70.3 71.1 54.5 62.4 STORM + T. Pooling 7B 25 1.82 128 256 71.3 72.5 59.5 63.4 STORM + T. Sampling * 7B 50 2.50 32 256 70.1 70.8 54.8 63.1 STORM + T. Pooling + T. Sampling * 7B 12.5 1.51 128 256 70.6 72.9 60.5 62.4 * 2 additional compression at test time.
@@
We evaluate our STORM across multiple configurations on recent long-video understanding benchmarks specifically designed to rigorously assess video-language model capabilities. These benchmarks include EgoSchema [39], MVBench [40], MLVU [41], LongVideoBench [42], and VideoMME [43]. We compare our approach against a wide range of representative video-language models, including recently proposed models tailored for long-video understanding [8, 21, 7, 20, 33, 15, 36]. See Table 1 for details.
@@ 4.2 Results on Video Understanding Benchmarks
We start with comparing our configuration STORM + T. Pooling (+ T. Sampling) with existing Video-LLMs. As shown in Table 1 and detailed in Table 2, STORM + T. Pooling achieves state-of-the-art performance across all long-video understanding benchmarks. Specifically, it achieves 71.3% accuracy on MVBench, 72.5% on MLVU, 59.5% on LongVideoBench, and 63.4% on VideoMME, outperforming all open-source Video-LLMs, including recent models specifically designed for long-context inputs such as LongVU and LongVILA. Additionally, our method significantly narrows the performance gap with proprietary models, outperforming GPT4-V and GPT4-O on MVBench and MLVU, as well as GPT4-V on VideoMME. Notably, STORM + T. Pooling achieves computational efficiency by compressing visual tokens to 25% of their original number before processing them through the LLM. We can further enhance efficiency by applying test-time temporal sampling in STORM + T. Pooling + T. Sampling while still achieving competitive results, reducing the token count to just 12.5% while maintaining competitive performance. In fact, this additional compression even improves results of certain benchmarks, yielding the best overall performance on MLVU and LongVideoBench.
@@
Next, we provide controlled comparisons within VILA-based models to reveal advantages of the Mamba-based temporal models in Table 2. We first compare the baseline VILA model with our STORM. By incorporating the Mamba module, STORM achieves performance gains on three out of four benchmarks, including a notable 2.3% improvement on VideoMME. Moreover, augmenting STORM with test-time temporal token sampling (STORM + T. Sampling) further enhances efficiency, reducing inference time by 43% while, surprisingly, maintaining or slightly boosting performance (an additional 0.8% gain on VideoMME). This advantageous behavior emerges because the Mamba module’s ability to effectively propagate temporal information across video frames, enabling redundant tokens to be discarded without compromising the model’s overall understanding.
@@ 5 Analysis
As illustrated in Figure 2, the Mamba-based temporal projector performs a spatio-temporal scan over input frames. This approach enables refinement of visual tokens before compression operations, thereby preserving key temporal and spatial cues—such as the positional information of frames being pooled—that would otherwise be lost in naive compression strategies. Consequently, the Mamba temporal module allows simple token compression methods to work more effectively for long videos compared to the baseline, as evidenced in Table 4 and further detailed in Table 9.
@@
Model Mamba T. T. S. Token Compression Latency (s) MVBench MLVU LongVideoBench VideoMME Module Sampling Pooling Pooling budget Ratio (%) w/o sub. VILA ✗ ✗ ✗ ✗ 8K 100 4.38 69.5 70.2 55.9 60.1 STORM ✓ ✗ ✗ ✗ 8K 100 4.26 70.3 71.1 54.5 62.4 ✓ ✓ ✗ ✗ 8K 50 2.50 70.1 70.8 54.8 63.1 ✓ ✗ ✓ ✗ 2K 25 1.82 70.4 71.0 54.2 61.2 ✓ ✗ ✗ ✓ 2K 25 1.82 68.9 69.2 56.0 61.1 ✓ ✓ ✓ ✗ 2K 12.5 1.51 70.1 71.0 54.0 60.8 ✓ ✓ ✗ ✓ 2K 12.5 1.51 68.9 69.5 56.3 60.9 ✓ ✗ ✓ ✓ 0.5K 6.25 1.36 68.5 68.2 53.7 60.2 ✓ ✓ ✓ ✓ 0.5K 3.13 1.29 67.2 68.6 55.3 60.1
@@
Our core design, temporal module + token compression, is model-agnostic and can be integrated into various video-LLM architectures. Table 6 shows the experiments on models utilizing different LLM models, vision encoders, and model sizes. Our design shows consistent improvements in both performance and latency for all configurations, clearly showcasing the universality of the model.
@@ 6 Conclusion
We introduced STORM, a novel Video-LLM architecture that enhances long-video understanding through a Mamba-based temporal encoder and efficient token reduction. By explicitly integrating spatiotemporal dynamics into visual tokens early in the pipeline, STORM enables significant token compression while preserving critical temporal information in the compressed inputs. Experiments demonstrate that STORM achieves new state-of-the-art results on long-video understanding benchmarks while substantially improving computational efficiency. Our work establishes a scalable approach to long video-language modeling that simultaneously achieve good performance and computation efficiency.

### v3

# STORM: Token-Efficient Long Video Understanding for Multimodal LLMs

@@ Abstract
Recent advances in video-based multimodal large language models (Video-LLMs) have significantly improved video understanding by processing videos as sequences of image frames. However, many existing methods treat frames independently in the vision backbone, lacking explicit temporal modeling, which limits their ability to capture dynamic patterns and efficiently handle long videos. To address these limitations, we introduce STORM (Spatiotemporal TOken Reduction for Multimodal LLMs), a novel architecture incorporating a dedicated temporal encoder between the image encoder and the LLM. Our temporal encoder leverages the Mamba State Space Model to integrate temporal information into image tokens, generating enriched representations that preserve inter-frame dynamics across the entire video sequence. This enriched encoding not only enhances video reasoning capabilities but also enables effective token reduction strategies, including test-time sampling and training-based temporal and spatial pooling, substantially reducing computational demands on the LLM without sacrificing key temporal information. By integrating these techniques, our approach simultaneously reduces training and inference latency while improving performance, enabling efficient and robust video understanding over extended temporal contexts. Extensive evaluations show that STORM achieves state-of-the-art results across various long video understanding benchmarks (more than 5% improvement on MLVU and LongVideoBench) while reducing the computation costs by up to and the decoding latency by 2.4-2.9 for the fixed numbers of input frames. Project website is available at here.
@@ 1 Introduction
Recent advancements in video-based multimodal Large Language Models (Video-LLMs) have improved AI systems’ ability to understand video content [1, 2, 3, 4, 5, 6]. These models typically process videos as sequences of individual frames, encoding each frame independently before feeding representations to a large language model (LLM) for temporal reasoning [3, 4, 2]. Despite leveraging powerful image encoders and LLMs [7, 8], these approaches exhibit fundamental limitations in video processing, particularly for long videos. The lack of explicit temporal encoding in image tokens forces LLMs to infer temporal relationships from static image sequences, creating substantial computational burden. Moreover, to manage computational overhead, existing methods like LongVILA [9] and LongVA [10] often use naive frame subsampling to reduce LLM processing tokens, resulting in significant information loss, potentially eliminating critical details necessary for comprehensive video understanding. Recently, LongVU [11] proposes token compression by adaptively selecting visual tokens based on user queries. However, unlike our approach, LongVU lacks a dedicated module for propagating spatiotemporal information before token selection, potentially limiting the model’s ability to preserve critical information after token reduction.
In this paper, we propose STORM (Spatiotemporal TOken Reduction for Multimodal LLMs), which introduces a novel temporal encoder between the image encoder and the LLM. This design integrates temporal dynamics earlier in the model pipeline, significantly enhancing the temporal reasoning capabilities of Video-LLMs while enabling substantial downstream token reduction. We employ Mamba [12] as the core of our temporal layer to allow both computation efficiency and generalization to extended temporal contexts. While there are existing Mamba-based video understanding models [13, 14, 15, 16, 17], unlike ours, these methods do not integrate Mamba with LLMs for multimodal tasks, and importantly, they mainly focus on replacing traditional backbones with Mamba architectures without explicitly leveraging Mamba’s ability for reducing video redundancy and enabling visual token compression.
@@
The key advantage of the Mamba layer is its ability to compress historical information into state representations. As consecutive frames in the video input often contain redundant information, our temporal encoder efficiently processes and propagates temporal information across the entire video sequence. The resulting visual tokens inherently encapsulate temporal history, allowing us to extract fewer tokens for LLM processing while preserving key information. We explore both training-free token subsampling and training-based compression methods (temporal and spatial token compression). Unlike previous frame-subsampling approaches, our method preserves essential temporal information in a compressed format, reducing computational load while improving performance through more comprehensive representation in a compact token space. Empirical evaluations demonstrate that our method significantly outperform baselines, especially on long-video inputs, while simultaneously achieving strong computational efficiency.
@@ 2 Related Work
Mamba for Video Understanding
Recent advances in linear state space models such as Mamba [12, 27] have sparked extensive exploration in applying them to video understanding tasks. Due to their sub-quadratic computation complexity, Mamba models achieve significant efficiency improvements compared to transformer-based architectures while still delivering competitive performance [12, 27, 28, 29, 30]. These properties make Mamba particularly suitable for video processing as the models are required to process long sequence inputs. For example, VideoMamba in [13] and VideoMamba (identical model naming) in [14] use Mamba-based visual backbone in video models and demonstrates the model’s strong ability to capture both local redundancy and long-term spatiotemporal dependencies. VideoMambaPro [15] proposes to improve Mamba’s video understanding ability by applying masking and residual connection during the backward scan. The Video Mamba Suite [16] further explores various architectures to integrate Mamba into existing video understanding models, demonstrating favorable efficiency-performance trade-offs for long sequence inputs. Mamba-ND [17] aims to improve Mamba’s performance on multi-dimensional data by investigating design choices such as SSM layer structure and scanning order within and across dimensions. However, unlike our approach, these works do not directly apply Mamba for Multimodal LLMs. More importantly, they primarily focus on replacing traditional backbones with Mamba architectures without explicitly leveraging Mamba’s unique ability to summarize historical information for reducing video redundancy and enabling visual token compression. Our paper addresses this research gap by proposing STORM, which proves to be both effective and efficient for video understanding while significantly reducing computational demands.
Concurrent Work
Recently, BIMBA [31] explores a similar architecture for long-video understanding, reporting similar benefits through empirical evaluation. We are encouraged to see these independent findings further support the hypothesis.
@@ 3 Method
This section introduces the Mamba-based temporal projector architecture and explores several token compression techniques for efficient long video processing. We investigate both training-based and training-free token compression strategies across temporal and spatial dimensions. We begin with an overview of our Mamba-based temporal projector, followed by a discussion of the token compression methods. The overview of our method is shown in Figure 2.
@@ 4 Experiments
4.1 Experiment Details
We use the pre-trained vision encoder in PaliGemma [37] and the LLM model from Qwen2-VL [38], and fine-tune them to adapt to our video datasets. The temporal projector is initialized with random weights. Each image is always resized to a resolution. Note that we also carried out experiments on more model configurations in Section 5 to showcase the universality of our design.
@@ 4.1 Experiment Details
VILA-Based Models Size Comp. Latency #Frames # Frames MVBench MLVU LongVideoBench VideoMME Ratio (%) (s) (train) (test) test dev val (w/o sub.) Duration 16 sec 3120 min 8 sec60 min 160 min Token Budget: 8K VILA Baseline 7B 100 4.31 32 256 69.5 70.2 55.9 60.1 STORM 7B 100 4.47 32 256 70.3 71.1 54.5 62.4 STORM + T. Pooling 7B 25 1.82 128 256 71.3 72.5 59.5 63.4 STORM + T. Sampling * 7B 50 2.50 32 256 70.1 70.8 54.8 63.1 STORM + T. Pooling + T. Sampling * 7B 12.5 1.51 128 256 70.6 72.9 60.5 62.4 * 2 additional compression at test time.
@@
We evaluate our STORM across multiple configurations on recent long-video understanding benchmarks specifically designed to rigorously assess video-language model capabilities. These benchmarks include MVBench [49], MLVU [50], LongVideoBench [51], and VideoMME [52]. We compare our approach against a wide range of representative video-language models, including recently proposed models tailored for long-video understanding [9, 11, 10, 26, 43, 21, 46]. See Table 1 for details.
@@ 4.2 Results on Video Understanding Benchmarks
We start with comparing our configuration STORM + T. Pooling (+ T. Sampling) with existing Video-LLMs. As shown in Table 1 and detailed in Table 2, STORM + T. Pooling achieves new state-of-the-art performance across all long-video understanding benchmarks. Specifically, it achieves 71.3% accuracy on MVBench, 72.5% on MLVU, 59.5% on LongVideoBench, and 63.4% on VideoMME, outperforming all open-source Video-LLMs, including recent models specifically designed for long-context inputs such as LongVU and LongVILA. Additionally, our method significantly narrows the performance gap with proprietary models, outperforming GPT4-V and GPT4-O on MVBench and MLVU, as well as GPT4-V on VideoMME. Notably, STORM + T. Pooling achieves computational efficiency by compressing visual tokens to 25% of their original number before processing them through the LLM. We can further enhance efficiency by applying test-time temporal sampling in STORM + T. Pooling + T. Sampling while still achieving competitive results, reducing the token count to just 12.5% while maintaining competitive performance. In fact, this additional compression even improves results of certain benchmarks, yielding the best overall performance on MLVU and LongVideoBench.
@@
Next, we provide controlled comparisons within VILA-based models to reveal advantages of the proposed Mamba-based temporal module in Table 2. We first compare the baseline VILA model with our STORM. By incorporating the Mamba module, STORM achieves performance gains on three out of four benchmarks, including a notable 2.3% improvement on VideoMME. Moreover, augmenting STORM with test-time temporal token sampling (STORM + T. Sampling) further enhances efficiency, reducing inference time by 43% while, surprisingly, maintaining or slightly boosting performance (an additional 0.8% gain on VideoMME). This advantageous behavior emerges because the Mamba module’s ability to effectively propagate temporal information across video frames, enabling redundant tokens to be discarded without compromising the model’s overall understanding.
@@ 5 Analysis
As illustrated in Figure 2, the Mamba-based temporal projector performs a spatiotemporal scan over input tokens. This enables refinement of visual tokens before compression operations, thereby preserving key temporal and spatial cues—such as the positional information of frames being pooled—that would otherwise be lost in naive compression strategies. Consequently, the Mamba temporal module allows simple token compression methods to work more effectively for long videos compared to the baseline, as evidenced in Table 4 and further detailed in Table 10.
@@
Model Mamba T. T. S. Token Compression Latency (s) MVBench MLVU LongVideoBench VideoMME Module Sampling Pooling Pooling budget Ratio (%) w/o sub. VILA ✗ ✗ ✗ ✗ 8K 100 4.31 69.5 70.2 55.9 60.1 STORM ✓ ✗ ✗ ✗ 8K 100 4.47 70.3 71.1 54.5 62.4 ✓ ✓ ✗ ✗ 8K 50 2.50 70.1 70.8 54.8 63.1 ✓ ✗ ✓ ✗ 2K 25 1.82 70.4 71.0 54.2 61.2 ✓ ✗ ✗ ✓ 2K 25 1.82 68.9 69.2 56.0 61.1 ✓ ✓ ✓ ✗ 2K 12.5 1.51 70.1 71.0 54.0 60.8 ✓ ✓ ✗ ✓ 2K 12.5 1.51 68.9 69.5 56.3 60.9 ✓ ✗ ✓ ✓ 0.5K 6.25 1.36 68.5 68.2 53.7 60.2 ✓ ✓ ✓ ✓ 0.5K 3.13 1.29 67.2 68.6 55.3 60.1
@@
Our core design, temporal module + token compression, is model-agnostic and can be integrated into various video-LLM architectures. Table 6 shows the experiments on models utilizing different LLM models (Qwen2 [38] vs Llama3 [53]), vision encoders (PaliGemma [37] vs SigLip [8]), model sizes (8B vs 8B vs 1.5B), and input resolutions ( vs ). Our design shows consistent improvements in both performance and latency for all configurations, clearly showcasing the universality of the model.
@@
Models 32F (T. Pooling) 128F (T. Pooling) VILA 58.9 61.7 Uni-dir STORM 62.2 62.5 Bi-dir STORM 61.2 63.4
Support for Streaming/Online Settings.
To evaluate the applicability of our method in streaming scenarios, we replaced the default bi-directional Mamba with a uni-directional variant, allowing the model to reuse prior states for constant-time computation as new frames arrive. The results are provided in Table 7. We find that both uni-directional and bi-directional variants significantly outperform the baseline without temporal modeling. Notably, the uni-directional Mamba performs competitively and even slightly surpasses the bi-directional counterpart when trained on 32-frame inputs. On the other hand, the bi-directional model demonstrates stronger performance as video length increases. These results highlight the critical role of the Mamba module and suggest its potential in enabling future designs that support streaming video input for video-LLMs.
@@ 6 Conclusion
We introduced STORM, a novel Video-LLM model that enhances long-video understanding using a Mamba-based temporal encoder and efficient token reduction. By explicitly integrating spatiotemporal dynamics into visual tokens early in the pipeline, STORM enables significant token compression while preserving critical information in the compressed inputs. Experiments demonstrate that STORM achieves new state-of-the-art results on long-video understanding benchmarks while substantially improving computational efficiency.
7 Acknowledgments
We would like to thank Xin Dong, Matthieu Le, Yunhao Fang, and Dacheng Li for their valuable discussions on this work. Sungjin Ahn was supported by the NRF of Korea funded by the Ministry of Science and ICT (MSIT) (No. 2021H1D3A2A03103645).